OpenStudy (anonymous):

**Multiple choice** The formula for the volume of a cone is V= volume of a cone. Find the radius, to the nearest hundredth, of a cone with a height of 5 in. and a volume of 20 in3. A. 3.82 in. B. 1.46 in. C. 1.95 in. D. 4.5 in.

OpenStudy (anonymous):



OpenStudy (anonymous):

r^2=3V/pie h

OpenStudy (anonymous):

The volume of a cone is \[V = \Pi r ^{2}h/ 3\] V= 20; h= 5 and Pi= 22/7 Pi*5* r^2/ 3 = 20 r^2 = 12/Pi r^2 = 3.81 Therefore r = Square root of 3.81 which approximately is 1.95. Hence, answer is 1.95 inches.